Solve the boundary value problem

y''+2y= −x

y(0)=0, y(1)+y'(1)=0


1
Expert's answer
2021-07-12T13:05:00-0400

The general solution can be written as "y=y_h+y_p."

Homogeneous differential equation


"y''+2y=0"

The characteristic polynomial 


"r^2+2=0"

"r_1=-i\sqrt{2}, r_2=i\sqrt{2}"

"y_h=c_1\cos (\sqrt{2}x)+c_2\sin (\sqrt{2}x)"

Let "y_p=Ax+B." Then "y_p''=0"


"2(Ax+B)=-x"

"A=-\dfrac{1}{2}, B=0"

"y_p=-\dfrac{1}{2}x"

"y=c_1\cos (\sqrt{2}x)+c_2\sin (\sqrt{2}x)-\dfrac{1}{2}x"

"y(0)=0:c_1\cos (\sqrt{2}(0))+c_2\sin (\sqrt{2}(0))-\dfrac{1}{2}(0)=0"

"c_1=0, y=c_2\sin (\sqrt{2}x)-\dfrac{1}{2}x"

"y'=c_2\cos (\sqrt{2}x)-\dfrac{1}{2}"

"y(1)+y'(1)=0"

"c_2\sin (\sqrt{2}(1))-\dfrac{1}{2}(1)+c_2\cos (\sqrt{2}(1))-\dfrac{1}{2}=0"

"c_2=\dfrac{1}{\sin(\sqrt{2})+\cos(\sqrt{2})}"

Then


"y=\dfrac{1}{\sin(\sqrt{2})+\cos(\sqrt{2})}\sin (\sqrt{2}x)-\dfrac{1}{2}x"